a health food store makes a blend of cereal by mixing oatmeal at $3.60 per kilogram and bran flakes at $4.80 per kilogram. how many kilograms of each should be mixed to obtain 30 kg of ceral worth $4.00 a kilogram?
Answer by stanbon(75887)

a health food store makes a blend of cereal by mixing oatmeal at $3.60 per kilogram and bran flakes at $4.80 per kilogram. how many kilograms of each should be mixed to obtain 30 kg of ceral worth $4.00 a kilogram?
-----------------------------------
Equation:
value + value = value
360x + 480(30 - x) = 400*30
360x + 480*30 - 480x = 400*30
-120x = -80*30
x = 20 kg (amt. of $3.60 oatmeal in the mixture)
-----------------------------
30-x = 10 kg (amt. of $4.80 bran flakes in the mixture)
====================================
Cheers,
Stan H.